Squareroot of 2x-5 + squareroot of 9+x=13
I square both sides and get 2x-5+9+x=169 and solve x=55 but it's wrong. What am I missing? Thanks for your help!

Answer by oscargut(891)

(a+b)^2=a^2+2ab+b^2
so
sqrt(2x-5)+sqrt(9+x)=13
2x-5+2sqrt(2x-5)sqrt(9+x)+9+x=169
3x+4+2sqrt(2x-5)sqrt(9+x)=169
3x-165 = -2sqrt(2x-5)sqrt(9+x)
(3x-165)^2=4(2x-5)(9+x)
if you need the rest please contact me